# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 2. A priori estimates and technical results
with $M>0$ to fix later. Then, we test (11) with $\rho_{\varepsilon,k+1}^{p-1}$ and we integrate in time in $(t_{1},t_{2})$, where $0\leq T_{k}\leq t_{1}\leq T_{k+1}\leq t_{2}\leq T$. We find
@@
$\frac{1}{p}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon,k+1}^{p})\,{\rm d}x\,{\rm d}t+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}C_{k+1}^{2}(p-1)\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p-2}\,{\rm d}x\,{\rm d}t.$ (15)
@@
$\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{2})+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}(p-1)\left(\frac{C_{k+1}}{C_{k+1}-C_{k}}\right)^{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t+\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{1}).$
@@
$\|\rho_{\varepsilon}\|_{L^{\infty}(0,T;\mathbb{R}^{d})}\leq M\sim C\left(1+\frac{1}{{\tau}}\right)^{\frac{d+2}{2p}}\|\rho_{\varepsilon}\|_{L^{p}(0,T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}}\left(1+\frac{1}{{\tau}^{\frac{d+2}{2p}}}\right)\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}.$
@@
3. Existence for the approximated equation
This section is devoted to show existence of solutions to (11) with smooth initial data. First, we semi-linearize (11) and obtain a monotone functional inside the divergence operator. Then, by standard theory for monotone operators we prove existence of this new problem. A Leray-Schauder fixed point will yield existence of solutions to (11).
3.1. Existence for the semi-linearized equation
For given $\varepsilon$, $\sigma\in[0,1]$ and $z\in L^{2}(0,T;L^{2}(\mathbb{R}^{d}))$, let us present the following regularized weak formulation of (9)
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\,\varphi\,{\rm d}x\,{\rm d}t=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
@@ 3.1. Existence for the semi-linearized equation
$\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$, where $\rho_{\varepsilon}(0,\cdot)=\rho_{{\rm in},\varepsilon}$ and $\rho_{{\rm in},\varepsilon}$ is a $W^{1,p}\cap L^{\infty}(\mathbb{R}^{d})$ approximation of $\rho_{{\rm in}}$ such that $\rho_{{\rm in},\varepsilon}\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ when $\varepsilon\rightarrow 0$. In order to prove the existence of the above equation, we use general theory of monotone operators, concretely, [22, Theorem 30.A]. We define the following operator $A$ as
$\langle A[u](t),\varphi\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u(t,x)\rvert}\right)_{+}\nabla u(t,x)\nabla\varphi\,{\rm d}x$
@@
1. Monotonicity. . The operator $A$ is monotone, since $\forall\;u,v\in H^{1}(\mathbb{R}^{d})$ we have
$\langle A[u]-A[v],u-v\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u-\left(1-\frac{\chi z_{+}}{\lvert\nabla v\rvert}\right)_{+}\nabla v\right]\nabla(u-v)\,{\rm d}x$
@@
$\lim_{t\rightarrow t_{0}}\langle A[u+tv],\varphi\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla(u+t_{0}v)\rvert}\right)_{+}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x+\varepsilon\int_{\mathbb{R}^{d}}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x$
@@
$\langle A[u],u\rangle=\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\lvert\nabla u\rvert^{2}\,{\rm d}x+\varepsilon\|u\|^{2}_{H^{1}(\mathbb{R}^{d})}\geq\varepsilon\|u\|_{H(\mathbb{R}^{d})}^{2},$
so $A$ is coercive.
@@
$=$ $\sup_{\begin{subarray}{c}\varphi\in H^{1}(\mathbb{R}^{d})\\ \|\varphi\|\leq 1\end{subarray}}\int_{\mathbb{R}^{d}}\left[\sigma\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u\nabla\varphi+\varepsilon\nabla u\nabla\varphi+\varepsilon u\varphi\right]\,{\rm d}x$
@@
Then, by [22, Theorem 30.A], there exists a unique weak solution $\rho_{\varepsilon}\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ to (3.1) with $\partial_{t}\rho\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$ satisfying the initial condition in the sense of the strong $L^{2}(\mathbb{R}^{d})$ convergence:
@@
$\mathcal{T}:(z,\sigma)\in L^{2}((0,T)\times\mathbb{R}^{d})\times[0,1]\mapsto\rho_{\varepsilon}\in L^{2}((0,T)\times\mathbb{R}^{d}),$
@@
3.2. Existence for approximated solutions
Now, we are going to use the Leray-Schauder fixed point theorem to show the existence of a solution to (11). Leray-Schauder theorem requires $\mathcal{T}$ to be compact. First, using $\rho_{\varepsilon}$ as test function in (3.1) we get
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon}^{2})\,{\rm d}x\,{\rm d}t=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(z_{+}+\varepsilon)}{\lvert\nabla\rho(t,x)\rvert}\right)_{+}|\nabla\rho_{\varepsilon}|^{2}\,{\rm d}x\,{\rm d}t$
@@ 3.2. Existence for approximated solutions
[…]
@@
$\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\phi_{R}\,{\rm d}x\leq$ $\;\;\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\,{\rm d}x+(\sigma+\varepsilon)\int_{0}^{T}\int_{B_{2R}}\frac{\rho_{\varepsilon}|\nabla\rho_{\varepsilon}|\cdot|x|}{\sqrt{1+\lvert x\rvert^{2}}}\,{\rm d}x\,{\rm d}t$
$+(\sigma+\varepsilon)\int_{0}^{T}\int_{B_{2R}}|\nabla\rho_{\varepsilon}|\rho_{\varepsilon}\sqrt{1+|x|^{2}}|\nabla\phi_{R}|\,{\rm d}x\,{\rm d}t.$
@@
Now, we have to prove that the map is continuous, i.e. that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\mathcal{T}(\sigma,z)$ when $z_{n}\rightarrow z$ strongly in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$ and $\sigma_{n}\rightarrow\sigma$ in $[0,1]$. Is easy to see that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma,z_{n})$. Then, as the map is compact, it only remains to prove
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}$ $(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)_{+}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t.$

### v3

@@ 2. A priori estimates
with $M>0$ to fix later. Then, we test (12) with $\rho_{\varepsilon,k+1}^{p-1}$ and we integrate in time in $(t_{1},t_{2})$, where $0\leq T_{k}\leq t_{1}\leq T_{k+1}\leq t_{2}\leq T$. We find that, using (14), we get
@@
$\frac{1}{p}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon,k+1}^{p})\,{\rm d}x\,{\rm d}t+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}C_{k+1}^{2}(p-1)\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p-2}\,{\rm d}x\,{\rm d}t.$ (17)
@@
$\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{2})+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}(p-1)\left(\frac{C_{k+1}}{C_{k+1}-C_{k}}\right)^{2}\int_{t_{1}}^{t_{2}}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t+\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{1}).$
@@
$\|\rho_{\varepsilon}\|_{L^{\infty}(0,T;\mathbb{R}^{d})}\leq M\sim C\left(1+\frac{1}{{\tau}}\right)^{\frac{d+2}{2p}}\|\rho_{\varepsilon}\|_{L^{p}(0,T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}}\left(1+\frac{1}{{\tau}^{\frac{d+2}{2p}}}\right)\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})}.$
We conclude this section by recalling a compactness criteria for functions defined in unbounded domains:
@@
3. Existence for the viscous problem
This section is devoted to the proof of existence of our equation equipped with extra regularizing terms $\varepsilon\Delta\rho+\varepsilon\rho$. More precisely, we will show existence of weak solution to
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\varphi dxdt=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt$ (19)
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\;dxdt,$
with initial data $\rho_{\varepsilon}(\cdot,0)=\rho_{in,\varepsilon}(\cdot)$, where $\rho_{in,\varepsilon}(\cdot)$ is an approximation of $\rho_{in}$ as smooth as one wishes.
The construction of solutions starts with a semi-linearization of (19). We replace $\left(1-\chi\frac{\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}$ with $\left(1-\chi\frac{z_{+}}{|\nabla\rho_{\varepsilon}|}\right)_{+}$ so that the diffusion term
$\left(1-\chi\frac{z_{+}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$
is monotone with respect to $\nabla\rho_{\varepsilon}$. Solution of this semilinearized problem follows from the theory of monotone operators (Section 3.1). Then, the Leray-Schauder fixed point theorem will yield existence of a solution to (19) (Section 3.2).
3.1. The semi-linearized equation
For given $\varepsilon$, $\tilde{\sigma}\in[0,1]$ and $z\in L^{2}(0,T;L^{2}(\mathbb{R}^{d}))$, let us consider
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\,\varphi\,{\rm d}x\,{\rm d}t=$ $-\tilde{\sigma}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
@@ 3.1. The semi-linearized equation
$\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$, where $\rho_{\varepsilon}(0,\cdot)=\rho_{{\rm in},\varepsilon}$ and $\rho_{{\rm in},\varepsilon}$ is a $W^{1,p}\cap L^{\infty}(\mathbb{R}^{d})$ approximation of $\rho_{{\rm in}}$ such that $\rho_{{\rm in},\varepsilon}\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ when $\varepsilon\rightarrow 0$.
We show the existence of solutions to (3.1) via the theory of monotone operators, concretely, [22, Theorem 30.A]. We define the following operator AA as
$\langle A[u](t),\varphi\rangle\;:=\;$ $\tilde{\sigma}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u(t,x)\rvert}\right)_{+}\nabla u(t,x)\nabla\varphi\,{\rm d}x$
@@
1. Monotonicity. . The operator AA is monotone, since $\forall\;u,v\in H^{1}(\mathbb{R}^{d})$ we have
$\langle A[u]-A[v],u-v\rangle=$ $\tilde{\sigma}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u-\left(1-\frac{\chi z_{+}}{\lvert\nabla v\rvert}\right)_{+}\nabla v\right]\nabla(u-v)\,{\rm d}x$
@@
$\lim_{t\rightarrow t_{0}}\langle A[u+tv],\varphi\rangle=$ $\tilde{\sigma}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla(u+t_{0}v)\rvert}\right)_{+}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x+\varepsilon\int_{\mathbb{R}^{d}}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x$
@@
$\langle A[u],u\rangle=\tilde{\sigma}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\lvert\nabla u\rvert^{2}\,{\rm d}x+\varepsilon\|u\|^{2}_{H^{1}(\mathbb{R}^{d})}\geq\varepsilon\|u\|_{H(\mathbb{R}^{d})}^{2},$
so AA is coercive.
@@
$=$ $\sup_{\begin{subarray}{c}\varphi\in H^{1}(\mathbb{R}^{d})\\ \|\varphi\|\leq 1\end{subarray}}\int_{\mathbb{R}^{d}}\left[\tilde{\sigma}\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u\nabla\varphi+\varepsilon\nabla u\nabla\varphi+\varepsilon u\varphi\right]\,{\rm d}x$
@@
Then, by [22, Theorem 30.A], there exists a unique weak solution $\rho_{\varepsilon}\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ to (3.1) with $\partial_{t}\rho_{\varepsilon}\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$ satisfying the initial condition in the sense of the strong $L^{2}(\mathbb{R}^{d})$ convergence:
@@
$\mathcal{T}:(z,\tilde{\sigma})\in L^{2}((0,T)\times\mathbb{R}^{d})\times[0,1]\mapsto\rho_{\varepsilon}\in L^{2}((0,T)\times\mathbb{R}^{d}),$
@@
3.2. The nonlinear problem
In this subsection we show that our map $\mathcal{T}$ satisfies the assumptions of the Leray-Schauder fixed point theorem. Namely we have to show that $\mathcal{T}$ is (i) continuous, (ii) compact, and (iii) constant when $\tilde{\sigma}=0$.
First, using $\rho_{\varepsilon}$ as test function in (3.1) we get
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon}^{2})\,{\rm d}x\,{\rm d}t=$ $-\tilde{\sigma}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(z_{+}+\varepsilon)}{\lvert\nabla\rho(t,x)\rvert}\right)_{+}|\nabla\rho_{\varepsilon}|^{2}\,{\rm d}x\,{\rm d}t$
@@ 3.2. The nonlinear problem
$\frac{1}{2}\|\rho_{\varepsilon}\|^{2}_{L^{\infty}(0,T;L^{2}(\mathbb{R}^{d}))}+\varepsilon\|\rho_{\varepsilon}\|^{2}_{L^{2}(0,T;H^{1}(\mathbb{R}^{d}))}\leq\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}^{2}\;dx.$
@@
$\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\phi_{R}\,{\rm d}x\leq$ $\;\;\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\,{\rm d}x+(\tilde{\sigma}+\varepsilon)\int_{0}^{T}\int_{B_{2R}}\frac{\rho_{\varepsilon}|\nabla\rho_{\varepsilon}|\cdot|x|}{\sqrt{1+\lvert x\rvert^{2}}}\,{\rm d}x\,{\rm d}t$
$+(\tilde{\sigma}+\varepsilon)\int_{0}^{T}\int_{B_{2R}}|\nabla\rho_{\varepsilon}|\rho_{\varepsilon}\sqrt{1+|x|^{2}}|\nabla\phi_{R}|\,{\rm d}x\,{\rm d}t.$
@@
Now, we have to prove that the map is continuous, i.e. that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\tilde{\sigma}_{n},z_{n})=\mathcal{T}(\tilde{\sigma},z)$ when $z_{n}\rightarrow z$ strongly in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$ and $\tilde{\sigma}_{n}\rightarrow\tilde{\sigma}$ in $[0,1]$. Is easy to see that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\tilde{\sigma}_{n},z_{n})=\lim\limits_{n\rightarrow\infty}\mathcal{T}(\tilde{\sigma},z_{n})$. Then, as the map is compact, it only remains to prove
@@
$\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\partial_{t}$ $(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\tilde{\sigma}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)_{+}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t.$